Destinywood:

Talib is trying to find the inverse of the function to the right. His work appears beneath it. Is his work correct? Explain your answer. f(x) = negative 8 x + 4. y = negative 8 x + 4. y minus 4 = negative 8 x. x = (y minus 4) divided by negative 8. f Subscript negative 1 Baseline (x) = (y minus 4) divided by negative 8.

jhonyy9:

do you think it in this way ???? f(x) = -8x +4 y = -8x +4 y -4 = -8x x = (y -4)/(-8) f^(-1) = (y -4)/(-8)